OpenStudy (anonymous):

someone please eplain why this is 18 ??? 3(2+4)

OpenStudy (anonymous):

i seems like 10?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

ohh nvm

OpenStudy (anonymous):

parenthasees

OpenStudy (anonymous):

first

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

yes exactly XD

ganeshie8 (ganeshie8):

Exactly !

OpenStudy (anonymous):

and this is what happens when you havent done math since highschool

ganeshie8 (ganeshie8):

PEMDAS and left to right precedence

OpenStudy (anonymous):

yea I know that

OpenStudy (anonymous):

good keep that in mind

OpenStudy (larseighner):

Because multiplication is distributive over addition 3(2 + 4) = 3(2) + 3(4) = 6 + 12 = 18 = 3(6) = 3(2 + 4)